# Conditional Density Estimations from Privacy-Protected Data

### v3

@@ Abstract
Many modern statistical analysis and machine learning applications require training models on sensitive user data. Differential privacy provides a formal guarantee that individual-level information about users does not leak. In this framework, randomized algorithms inject calibrated noise into the confidential data, resulting in privacy-protected datasets or queries. However, restricting access to only privatized data during statistical analysis makes it computationally challenging to make valid inferences on the parameters underlying the confidential data. In this work, we propose simulation-based inference methods from privacy-protected datasets. In addition to sequential Monte Carlo approximate Bayesian computation, we use neural conditional density estimators as a flexible family of distributions to approximate the posterior distribution of model parameters given the observed private query results. We illustrate our methods on discrete time-series data under an infectious disease model and with ordinary linear regression models. Illustrating the privacy-utility trade-off, our experiments and analysis demonstrate the necessity and feasibility of designing valid statistical inference procedures to correct for biases introduced by the privacy-protection mechanisms.
@@ 1 Introduction
Motivation. Many AI systems require collecting and training on massive amounts of personal information (such as income, disease status, location, purchase history, etc). Despite unprecedented data collection efforts by companies, governments, researchers, and other agencies, oftentimes the data collectors have to lock the data inside their own database due to privacy concerns. Differential privacy (DP) provides a mathematical definition for privacy protection of individual data. Under this framework, privacy-protecting procedures (i.e., DP algorithms) have enabled data collectors, such as tech companies, the US Census Bureau, and social scientists, to share research data in a wide variety of settings while protecting the privacy of individual users. Privacy researchers typically collect confidential data and then inject calibrated random noise into the confidential data to achieve the desired levels of privacy protection. Some algorithms aim for DP data analysis, resulting in DP optimizations (Arora et al.,, 2023; Bassily et al.,, 2021), approximations (Chaudhuri et al.,, 2013; Bie et al.,, 2023), or predictions (Rho et al.,, 2023); while other algorithms produce DP datasets (or descriptive statistics), which enables data sharing across research teams and entities. Examples of the latter include the 2020 US Census (Abowd et al.,, 2022; Gong et al.,, 2022; Drechsler,, 2023) and Facebook URL dataset (Evans and King,, 2023). Our work focuses on this data sharing regime and complex models (e.g. continuous-time Markov jump processes).
@@
While the DP data sharing regime corrupts confidential information in order to satisfy privacy, since the probabilistic design of such mechanisms can be publicly known, in principle, analysts can still conduct reliable estimation and uncertainty quantification by accounting for bias and noises introduced during privatization. In practice, however, valid inference based on privatized data is a challenge that requires the revision of existing statistical methods designed originally for confidential data (Foulds et al.,, 2016). Even for well-understood procedures such as ordinary linear regression and generalized linear models, adding the extra layer of privacy protection has introduced new theoretical and methodological questions in statistics (Cai et al.,, 2021; Alabi and Vadhan,, 2022; Li et al.,, 2023; Barrientos et al.,, 2019)
@@
In many applications, even the confidential data likelihood functions are intractable or time-consuming to compute. Accounting for privacy noise on top of model complexity is a formidable challenge. Without developing valid inference procedures under this regime, we must either make biased estimations or restrict ourselves to simple models. In this work, we propose methods to estimate the parameters of complex models that underlie privacy-protected data.
@@
Related works. There is a fast-growing literature on statistical inference under differential privacy. We point out several Bayesian inference methods from privatized data. Markov Chain Monte Carlo (MCMC) methods have been proposed in some specific models and priors, such as exponential family distributions (Bernstein and Sheldon,, 2018) and Bayesian linear regression (Bernstein and Sheldon,, 2019). As for generic algorithms, Ju et al., (2022) proposed a data-augmentation MCMC strategy to overcome the intractable marginal likelihood resulting from privatization, and Gong, (2022) derived point estimates of the posterior distribution using the Expectation-Maximization algorithm. Several frequentist inference methods have also been developed. Karwa et al., (2015) employed a parametric bootstrap method to construct confidence intervals for the model parameters of the log-linear model. Awan and Wang, (2023) proposes simulation-based inference methods for hypothesis testing and confidence intervals.
@@
To the best of our knowledge, only two papers (Waites and Cummings,, 2021; Su et al.,, 2023) have incorporated normalizing flows (Kobyzev et al.,, 2020; Papamakarios et al.,, 2021) and DP, and both works design DP-versions of normalizing flow. In contrast, our work uses flow-based methods as a neural density estimation tool to analyze DP-protected data.
@@
Our contributions. In this work, we propose several likelihood-free inference methods that make statistical inferences from privacy-protected data. In addition to sequential Monte Carlo approximate Bayesian computation (SMC-ABC), we use sequential neural density estimation methods, which use neural networks as a flexible family of distributions, to approximate the private data posterior distribution. Unlike existing methods for private data, these simulation-based methods require only simulations from a generative model for confidential data. We demonstrate the efficiency and utility of our methods on an infectious disease model using synthetic and real disease outbreak data. We also propose a privacy mechanism for the release of the infection curve. Our experiments also demonstrate the privacy-utility trade-off in linear regression.
@@ 2 Background and challenges
Let $\theta\in\Theta$ be the model parameter and $x=(x_{1},\cdots,x_{n})\in\mathbb{X}^{n}$ represent the confidential database, containing a total of $n$ records. We model the database with some likelihood function $f(x\mid\theta)$. This confidential data model can be a ‘simulator’ whose likelihood can be impossible to compute. Before adding privacy, suppose that we observe some sample $x^{o}$; our inference objective is to perform inference on the posterior distribution $\pi(\theta\mid x^{o})\propto\pi(\theta)f(x^{o}\mid\theta)$ where $\pi(\theta)$ is the prior. Our work studies posterior inference under the constraint of DP, where we must learn about $x^{o}$ through some DP query result $s_{\mathrm{dp}}$ instead of through $x^{o}.$
Differential privacy. Given confidential data $x$, let $\eta$ be a randomized algorithm to produce a ‘differentially private statistic’ $s_{\mathrm{dp}}$ from $x$. We also use $\eta(s_{\mathrm{dp}}\mid x)$ to denote the conditional density of the private output $s_{\mathrm{dp}}$ given the confidential data $x$.
@@
Intuitively, privacy protection is achieved when perturbing one individual’s response in the dataset leads to only a small change on the algorithm’s outcome. We present the definition of $\epsilon$-DP from Dwork et al., (2006), where neighboring databases induce ‘close’ distributions on $s_{\mathrm{dp}}.$
@@
A randomized algorithm $\eta$ satisfies $\epsilon$-DP if for all possible values of $s_{\mathrm{dp}}$ and for all pairs of ‘neighboring’ databases $(x,x^{\prime})\in\mathbb{X}^{n}\times\mathbb{X}^{n}$, which are databases differing by only one record [denoted by $d(x,x^{\prime})\leq 1$], the following probability ratio is bounded:
$\frac{\int_{A}\eta(s_{\mathrm{dp}}\mid x)\ \mathrm{d}s_{\mathrm{dp}}}{\int_{A}\eta(s_{\mathrm{dp}}\mid x^{\prime})\ \mathrm{d}s_{\mathrm{dp}}}\leq\exp(\epsilon),\quad\forall A\subseteq\mathrm{Image}(\eta).$ (1)
@@
A popular class of DP mechanism is output perturbation, where privacy is achieved by first computing a query $s:\mathbb{X}^{n}\to\mathbb{S}$ (such as mean, median, histogram, contingency tables) of the database and then releasing $s(x)$ with some noise. Since many queries are embedded in lower dimensional spaces than the confidential data, this might facilitate efficient computations. In Section 3.3, we leverage the low dimensionality of $s_{\mathrm{dp}}$ by incorporating quasi Monte Carlo techniques.
To satisfy $\epsilon$-DP, the query result $s$ must have finite sensitivity.
@@
The Laplace mechanism is a popular privacy protection technique.
@@
Intractable likelihood and posterior. A key challenge with data analysis on privatized data is the intractable likelihood.
@@
The data augmentation strategy circumvents the intractability of Eq.(2) by working with the joint posterior distribution $\pi(\theta,x\mid s_{\mathrm{dp}})\propto f(x\mid\theta)\pi(\theta)\eta(s_{\mathrm{dp}}\mid x)$ instead of the marginal posterior in Eq.(3).
In this work, we study the challenging scenarios when the confidential likelihood $f(x\mid\theta)$ is no longer analytically tractable or is time-consuming to compute. This situation is ‘triply intractable’, as there are three levels of intractability in $f(x\mid\theta)$, $f(s_{\mathrm{dp}}\mid\theta)$, and $\pi(\theta\mid s_{\mathrm{dp}}).$
Likelihood-free inference. For complex confidential data generating processes (Gourieroux et al.,, 1993; Brehmer et al.,, 2020), several studies have trained conditional density estimators to perform likelihood-free inference. Recent advances have embraced the use of neural networks, in particular, normalizing flows (Dinh et al.,, 2017; Papamakarios et al.,, 2017; Durkan et al.,, 2019; Papamakarios et al.,, 2021) as a highly flexible family of conditional densities.
We review the idea of neural density approximations on the confidential data posterior. Using the variational family $\{q_{\phi}(\theta\mid x)\}$ to approximate a target posterior distribution $\pi(\theta\mid x^{o})\propto f(x^{o}\mid\theta)\pi(\theta)$, we can set up a loss function $\ell(\phi)=-\sum_{i=1}^{N}\log q_{\phi}(\theta^{(i)}\mid x^{(i)})$, where the training data $\{(\theta^{(i)},x^{(i)})\}_{i=1}^{N}$ are sampled from the joint probability density $p(\theta,x)=\pi(\theta)f(x\mid\theta)$. The loss function $\ell(\phi)$ is a sample-based approximation of $\mathbb{E}_{p(\theta,x)}\left[-\log q_{\phi}(\theta\mid x)\right]$. Recognizing that it can be inefficient to approximate the above integral with samples from the joint model $p(\theta,x)$, we can design sequential training procedures (Papamakarios and Murray,, 2016; Lueckmann et al.,, 2017; Greenberg et al.,, 2019) that gradually improve the proposal distribution by using the current approximation as the proposal distribution for the next round. Other neural density estimation approaches include approximating likelihood functions (Papamakarios et al.,, 2019) or likelihood ratios (Miller et al.,, 2022). We refer the readers to Cranmer et al., (2020); Lueckmann et al., (2021) for systematic reviews on the topic. Finally, simulation-based inference under model misspecification has recently been studied in Ward et al., (2022) and Kelly et al., (2023).
3 Methods
We aim to train conditional density estimators to estimate the private data posterior $\pi(\theta\mid s_{\mathrm{dp}})$. We consider two complementary approaches: 1) approximate the private data marginal likelihood $f(s_{\mathrm{dp}}\mid\theta)$ and 2) approximate $\pi(\theta\mid s_{\mathrm{dp}})$ directly.

### v4

# Simulation-based Bayesian Inference from Privacy Protected Data

@@ Abstract
Many modern statistical analysis and machine learning applications require training models on sensitive user data. Under a formal definition of privacy protection, differentially private algorithms inject calibrated noise into the confidential data or during the data analysis process to produce privacy-protected datasets or queries. However, restricting access to only privatized data during statistical analysis makes it computationally challenging to make valid statistical inferences. In this work, we propose simulation-based inference methods from privacy-protected datasets. In addition to sequential Monte Carlo approximate Bayesian computation, we adopt neural conditional density estimators as a flexible family of distributions to approximate the posterior distribution of model parameters given the observed private query results. We illustrate our methods on discrete time-series data under an infectious disease model and with ordinary linear regression models. Illustrating the privacy-utility trade-off, our experiments and analysis demonstrate the necessity and feasibility of designing valid statistical inference procedures to correct for biases introduced by the privacy-protection mechanisms.
@@ 1 Introduction
Many AI systems require collecting and training on massive amounts of personal information (such as income, disease status, location, purchase history, etc.). Despite unprecedented data collection efforts by companies, governments, researchers, and other agencies, oftentimes, data collectors have to lock the data inside their own databases due to privacy concerns. Differential privacy (DP) provides a mathematical definition for the protection of individual data (Drechsler et al.,, 2024). Under this framework, privacy-protecting procedures (i.e., DP algorithms) have enabled data collectors, such as tech companies, the US Census Bureau, and social scientists, to share research data in a wide variety of settings while protecting the privacy of individual users. Privacy researchers typically collect confidential data and then inject calibrated random noise into the confidential data to achieve the desired levels of privacy protection. Some algorithms aim for DP data analysis, resulting in DP optimizations (Arora et al.,, 2023; Bassily et al.,, 2021), approximations (Chaudhuri et al.,, 2013; Bie et al.,, 2023), or predictions (Rho et al.,, 2023); while other algorithms produce DP datasets (or descriptive statistics), which enables data sharing across research teams and entities. Examples of the latter include the 2020 US Census (Abowd et al.,, 2022; Gong et al.,, 2022; Drechsler,, 2023), the Facebook URL dataset (Evans and King,, 2023), and New York Airbnb Open Data (Guo and Hu,, 2023). Our work tackles this data-sharing regime: we aim to make valid statistical inferences with privatized data, and we place a special focus on using complex models such as continuous-time Markov jump processes.
@@
Although the DP data-sharing regime corrupts confidential information in order to satisfy privacy, since the probabilistic design of such mechanisms can be publicly known, in principle analysts can still conduct reliable estimation and uncertainty quantification by accounting for bias and noises introduced during privatization. However, in practice, valid inference based on privatized data is a challenge that requires the revision of existing statistical methods designed originally for confidential data (Foulds et al.,, 2016). Even for well-understood procedures such as ordinary linear regression and generalized linear models, adding the extra layer of privacy protection has introduced new theoretical and methodological questions in statistics (Cai et al.,, 2021; Alabi and Vadhan,, 2022; Li et al.,, 2023; Barrientos et al.,, 2019).
@@
However, for complex models, even the confidential data likelihood functions are intractable or time-consuming to compute. Accounting for privacy noise on top of that is a formidable challenge. Without developing valid inference procedures under this regime, we must either make biased estimations or restrict ourselves to simple models. In this work, we propose methods to estimate the parameters of complex models that underlie privacy-protected data.
Related works.
@@
There is a fast-growing literature on statistical inference under differential privacy. Early work by Williams and McSherry, (2010) first proposed using Bayesian inference to handle DP noise. Since then, various Bayesian methods have been developed for privatized data. Markov Chain Monte Carlo (MCMC) methods have been proposed in some specific models and priors, such as exponential family distributions (Bernstein and Sheldon,, 2018), Bayesian linear regression (Bernstein and Sheldon,, 2019) and generalized linear models (Kulkarni et al.,, 2021). As for generic algorithms, Ju et al., (2022) proposed a data-augmentation MCMC strategy to overcome the intractable marginal likelihood resulting from privatization, and Gong, (2022) derived point estimates of the posterior distribution using the Expectation-Maximization algorithm. Several frequentist inference methods have also been developed. Karwa et al., (2015) employed a parametric bootstrap method to construct confidence intervals for the model parameters of the log-linear model. Awan and Wang, (2023) proposes simulation-based inference methods for hypothesis testing and confidence intervals, for frequentist inference.
@@
To the best of our knowledge, only three papers (Waites and Cummings,, 2021; Lee et al.,, 2022; Su et al.,, 2023) have incorporated normalizing flows (Kobyzev et al.,, 2020; Papamakarios et al.,, 2021) and DP, and both works design DP-versions of normalizing flow. In contrast, our work uses flow-based methods as a neural density estimation tool to analyze DP-protected data.
Our contributions.
@@
In this work, we propose several likelihood-free inference methods that make statistical inferences from privacy-protected data. First, we highlight that sequential Monte Carlo approximate Bayesian computation (SMC-ABC) can be used for this purpose, which improves the current practice of using ABC (Gong,, 2022). Next, we propose sequential private posterior estimation (SPPE) and sequential private likelihood estimation (SPLE), two sequential neural density estimation methods to approximate the private data posterior distribution with normalizing flows. Unlike likelihood-based methods to learn from private data, SMC-ABC, SPPE, and SPLE require only simulations from a generative model for confidential data, and hence are applicable to more models. We demonstrate the efficiency of our methods on an infectious disease model using synthetic and several real disease outbreak data. SPPE and SPLE require fewer numbers of simulations from the data model to achieve the same inference results compared with SMC-ABC. We also propose a privacy mechanism for the release of the infection curve. Our experiments also demonstrate the privacy-utility trade-off in linear regression, where we compared the proposed likelihood-free methods with data augmentation MCMC, a likelihood-based inference method. The code is available on GitHub [^1] https://github.com/Yifei-Xiong/Simulation-based-Bayesian-Inference-from-Privacy-Protected-Data.
@@ 2 Background and challenges
Let $\theta\in\Theta$ be the model parameter and $x=(x_{1},\cdots,x_{n})\in\mathbb{X}^{n}$ represent the confidential database, where $n$ is the sample size. We model the database with some likelihood function $f(x\mid\theta)$. This confidential data model can be a ‘simulator’ whose likelihood can be impossible to compute. In the confidential data setting, our interest is to approximate the posterior distribution $\pi(\theta\mid x^{o})\propto\pi(\theta)f(x^{o}\mid\theta)$ where $x^{o}$ is the observed dataset and $\pi(\theta)$ is the prior. In this section, we first review some key approaches for likelihood-free inference in the confidential data setting and then describe output perturbation mechanisms to generate differentially private queries $s_{\mathrm{dp}}$. Finally, we explain why learning $\theta$ given observed $s_{\mathrm{dp}}^{o}$ is challenging. We will present new methods for likelihood-free inference in the privatized data setting in Section 3.
2.1 Likelihood-free inference for confidential data
Likelihood-free inference methods require only the ability to generate data from a simulator model, instead of evaluating its likelihood.
Approximate Bayesian Computation (ABC) and sequential Monte Carlo ABC (SMC-ABC).
ABC-based methods circumvent likelihood evaluations with simulations. In its most basic form, a set of parameters are sampled from the prior $\theta^{(i)}\sim\pi(\theta)$ and a data set $x^{(i)}$ is simulated from the model for each $\theta^{(i)}$. The samples $\theta^{(i)}$ such that $x^{(i)}$ is close to the observed data $x^{o}$ provide an approximation to the posterior $\pi(\theta\mid x^{o}).$ It is also common to choose some summary statistics $s$ and to retain samples according to distance between each $s(x^{(i)})$ and $s(x^{o})$. The efficiency of ABC is controlled by the closeness between the prior distribution, which can be viewed as its proposal distribution, and the posterior distribution, which is the target distribution. SMC-ABC (Sisson et al.,, 2007; Beaumont et al.,, 2009) improves ABC by specifying a sequence of intermediate target distributions between the prior and the posterior, and thus the $\theta^{(i)}$ samples can gradually evolve towards the target distribution. The sequential updates in SMC-ABC propagate and reweight the parameter samples with importance sampling.
Neural density estimation.
In contrast to sampling-based approximations like ABC and SMC-ABC, we can pursue density approximation of $\pi(\theta\mid x^{o})$ with neural networks. We illustrate a neural estimator for the posterior density $\pi(\theta\mid x)$. Let $\{q_{\phi}(\theta\mid x)\}_{\phi}$ denote the collection of densities representable by an architecture $q$ and weight parameters $\phi$. We can train the neural estimator to minimize an ideal loss function
$\hat{\phi}=\arg\min_{\phi}\mathbb{E}_{p(\theta,x)}\left[-\log q_{\phi}(\theta\mid x)\right]=\arg\min_{\phi}\int_{\Theta\times\mathbb{X}^{n}}\left[-\log q_{\phi}(\theta\mid x)\right]\cdot\pi(\theta)f(x\mid\theta)\mathrm{d}\theta\mathrm{d}x.$
This loss is an expected value of $-\log q_{\phi}$ under the joint distribution $p(\theta,x)=\pi(\theta)f(x\mid\theta)$. When $f(x\mid\theta)$ is intractable, the expectation is also intractable, and training relies on Monte Carlo approximations of the integral using samples $\{(\theta^{(j)},x^{(j)})\}_{i=1}^{N}$. In sequential neural density estimation, we can sequentially improve the Monte Carlo estimation of the loss function (Papamakarios and Murray,, 2016; Lueckmann et al.,, 2017; Greenberg et al.,, 2019). Later, in Section 3.4, we will use randomized quasi-Monte Carlo (RQMC) methods (Owen, 1997a, ; Owen, 1997b, ) as a drop-in replacement for standard Monte Carlo to reduce variance. Other neural density approaches include approximating likelihood functions (Papamakarios et al.,, 2019) or likelihood ratios (Miller et al.,, 2022). We refer the readers to (Cranmer et al.,, 2020; Lueckmann et al.,, 2021) for systematic reviews of neural density approximations.
A notable neural density estimation method is normalizing flows (NF) (Dinh et al.,, 2017; Papamakarios et al.,, 2017; Durkan et al.,, 2019; Papamakarios et al.,, 2021), which are adopted in our experiments. NF start with a simple base density (e.g. multivariate Normal) and push it through a sequence of invertible and differentiable transformations. The resulting distribution is richer and more complex. We choose NF because of their expressive power and popularity in probabilistic modeling.
2.2 Differential privacy
@@
Given confidential data $x$, let $\eta$ be a randomized algorithm to produce a ‘differentially private statistic’ $s_{\mathrm{dp}}$ from $x$. We also use $\eta(s_{\mathrm{dp}}\mid x)$ to denote the conditional density of the private output $s_{\mathrm{dp}}$ given the confidential data $x$. Intuitively, a procedure is private when perturbing one individual’s response in the dataset leads to only a small change in the algorithm’s outcome. This is characterized by the $\epsilon$-DP definition from Dwork et al., (2006), based on neighboring databases.
@@
We say $x,x^{\prime}\in\mathbb{X}^{n}$ are ‘neighboring’ databases if they differ by one and only one data record. Denote this by $d(x,x^{\prime})\leq 1$. A privacy mechanism with conditional density $\eta$ satisfies $\epsilon$-DP if, for all possible values of $s_{\mathrm{dp}}$ and for neighboring datasets, the following probability ratio is bounded:
$\frac{\int_{A}\eta(s_{\mathrm{dp}}\mid x)\ \mathrm{d}s_{\mathrm{dp}}}{\int_{A}\eta(s_{\mathrm{dp}}\mid x^{\prime})\ \mathrm{d}s_{\mathrm{dp}}}\leq\exp(\epsilon),\quad\forall A\subseteq\mathrm{Range}(\eta).$ (1)
@@
Output perturbation methods achieve privacy by first computing a query $s:\mathbb{X}^{n}\to\mathbb{S}$ (such as mean, median, histogram, contingency tables) of the database and then releasing $s(x)$ with added noise. To satisfy $\epsilon$-DP, the query $s$ must have finite sensitivity.
@@
A common output perturbation technique is the Laplace mechanism.
@@
We exploit output perturbation on summary statistics in two ways. From an inference perspective, when $s(x)$ is a sufficient statistic for $\theta$, the confidential data posterior $\pi(\theta\mid x)$ is fully characterized by $s(x)$ and the prior. This is the rationale for choosing $\pi(\theta\mid s_{\mathrm{dp}})$ as the target distribution and for the posterior from perturbed data to be still informative about the model parameters. From a computation perspective, since many summary statistics are of lower dimension than the confidential data, this can facilitate efficient computations. In Section 3.4, we leverage the low dimensionality of $s_{\mathrm{dp}}$ by incorporating quasi-Monte Carlo techniques.
2.3 Bayesian inference with privatized queries
In the private data setting, our goal is to approximate the posterior distribution of $\theta$ given privatized data $s_{\mathrm{dp}}.$ A key challenge with data analysis on privatized data via $\pi(\theta\mid s_{\mathrm{dp}})$ is the intractable private data marginal likelihood.
@@ 2.3 Bayesian inference with privatized queries
The data augmentation strategy circumvents the intractability of evaluating Equation (2) by working with the joint posterior distribution $p(\theta,x\mid s_{\mathrm{dp}})\propto f(x\mid\theta)\pi(\theta)\eta(s_{\mathrm{dp}}\mid x)$ instead of the marginal posterior in Equation (3).
In this work, we study the challenging scenarios when the confidential likelihood $f(x\mid\theta)$ is intractable. This situation is ‘triply intractable’, as there are three levels of intractability in $f(x\mid\theta)$, $f(s_{\mathrm{dp}}\mid\theta)$, and $\pi(\theta\mid s_{\mathrm{dp}}).$ Data augmentation MCMC, which includes an inference step to sample from $\pi(\theta\mid x)$, is no longer applicable in the ‘triply intractable’ setting.
3 Proposed methods
We present methods for likelihood-free inference for the triply intractable posterior distribution $\pi(\theta\mid s_{\mathrm{dp}}).$ First and foremost, we recognize that SMC-ABC methods are viable solutions to approximate the private data posterior, as long as the privacy mechanism is publicly known and can be replicated by the data analyst.
Next, under the same assumptions, we focus on two neural density approximation methods that leverage state-of-the-art simulation-based inference methods and can be more efficient than the SMC-ABC baseline. We present two complementary approaches: a) Section 3.2: approximate the private data marginal likelihood $f(s_{\mathrm{dp}}\mid\theta)$ in Equation (2), and b) Section 3.3: approximate $\pi(\theta\mid s_{\mathrm{dp}})$ from Equation (3) directly.
3.1 Adapting SMC-ABC to the private data setting
A data analyst can create a (hierarchical) private data simulator for $s_{\mathrm{dp}}$ by combining the confidential data simulator $x\sim f(\cdot\mid\theta)$ and a tractable output perturbation mechanism (e.g. the Laplace mechanism) with $s_{\mathrm{dp}}\sim\eta(\cdot\mid x)$. With this simulator for $s_{\mathrm{dp}}$, a rejection ABC algorithm can simulate parameter samples $\{\theta^{(i)}\}_{i=1,\ldots,N}$ from the prior, and then keep samples correspond to a simulated private query $s_{\mathrm{dp}}^{(i)}$ close to the observed value $s_{\mathrm{dp}}^{o}$. This insight has been exploited in Gong, (2022), using ABC to approximate $\pi(\theta\mid s_{\mathrm{dp}}^{o}$). We point out that SMC-ABC can also be used for this purpose. We include a description of the algorithm in Appendix C for completeness.
Both SMC-ABC and ABC aim to target the exact posterior distribution, unlike neural density estimation methods which insist that the approximation come from some parametric family $q_{\phi}$. For this reason, we will use SMC-ABC as a baseline to test and validate our methods in Section 4.